OpenStudy (anonymous):

Solve the following system of equations: x + y − 4z = −7 2x + 3y − 3z = −6 2x + 2y − z = 0

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

do you know how to solve "system of equations" with 2 variables?

OpenStudy (anonymous):

yah substitute but I don't know how to do it with 3

OpenStudy (jdoe0001):

have you done solving by elimination yet?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

lemme use the elimination... one sec

OpenStudy (jdoe0001):

the way it works is, we pick 2 off the set of 3 then we eliminate one of the variables then we pick another 2 off the set, and do the same so let me pick these two for now \(\bf {\color{brown}{ bx + y - 4z = -7 \\ 2x + 3y - 3z = -6}}\\ 2x + 2y - z = 0 \) \(\large \begin{array}{llll} x + y - 4z = -7 & {\color{red}{ \times -3}}\to &-3x-3y\cancel{ +12z }=+21\\ 2x + 3y - 3z = -6&{\color{red}{ \times 4}}& 8x+12y\cancel{ -12z }=-24\\ \hline\\ &&{\color{blue}{ 5x+9y=\qquad -3}} \end{array}\)

OpenStudy (anonymous):

ok then?

OpenStudy (anonymous):

ok then the last 2

OpenStudy (jdoe0001):

then you grab the "resultant 2 equations" and solve for either "x" or "y" to get the other \(\bf {\color{blue}{ 5x+9y=-3\\ 9x-7y=-7}}\)

OpenStudy (anonymous):

wait im confused -.- could you show me or tell me what to do? like add them or get x or y by itself to substitute?

OpenStudy (jdoe0001):

hmmm actually I missed something.... the -4 should give use -8x

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} x + y - 4z = -7 && x + y \cancel{ - 4z } = -7\\ 2x + 2y - z = 0 &{\color{red}{ \times -4}}\to &-8x-8y\cancel{ +4z }=0\\ \hline\\ &&{\color{blue}{ -7x-7y=\qquad -7}} \end{array}\)

OpenStudy (jdoe0001):

well notice how I added those 2 there is just a straight forward vertically, once you have the "x" and "y" an "z" all lined up vertically then you add them vertically :)

OpenStudy (anonymous):

so then how do you solve?

OpenStudy (jdoe0001):

so the 2 resultant equation will be \(\bf 5x+9y=-3\\ -7x-7y=-7\)

OpenStudy (jdoe0001):

well use the elimination method

OpenStudy (jdoe0001):

just find 1 of the variables firstly, say "x" or "y"

OpenStudy (anonymous):

so 2x+2y=-10 which would be lets say y=-x-5?

OpenStudy (jdoe0001):

using substitution, yes

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

where did we get 2x+2y=-10 again?

OpenStudy (anonymous):

^^^ from subtracting the last 2 equations you posted

OpenStudy (jdoe0001):

the ones that you'd need to solve will be the "resulting equations" from the elimination or 5x + 9y = -3 -7x -7y = -7

OpenStudy (jdoe0001):

ohh .. I see

OpenStudy (jdoe0001):

yea.. that equation should also be workable to find "x" or "y"

OpenStudy (jdoe0001):

so if y = -x-5 then substituting that on either equation say the 5x+9y=-3 then \(\bf {\color{brown}{ y}}=-x-5\qquad 5x+9{\color{brown}{ y}}=-3 \to 5x+9{\color{brown}{ (-x-5)}}=-3\)

OpenStudy (anonymous):

so -5x^2-25x-9x-45=-3?

OpenStudy (anonymous):

I honestly don't know like how to solve this an its due at 6:30 :(

OpenStudy (anonymous):

and I just have 2 left... do you think we can do the second one?

OpenStudy (jdoe0001):

well... you can always post anew... .if I dunno someone else may, and we can also revise each other